You believe you make money ‘arbitraging’ differences between retailer’s sales prices by buying at local stores and reselling on Amazon marketplace. You need to invest $100,000 today to get the project going today for a warehouse, servers, and vehicles. You expect you can earn $25,000 per year for the next 5 years starting a year from today. At the end of the project, you expect to pay out $50,000 to cover lawsuits and seller disputes. If your discount rate is 4%, what is the IRR and NPV and should you take the project?
Answer #1

NPV =PV of Income -PV of Lawsuit -Initial Investment =25000*((1-(1+4%)^-5)/4%)-50000/(1+5%)^5-100000 =-27880.75

Cash flow 5 =25000-50000 =-25000
IRR using financial calculator
CF0=-100000;CF1 =25000;CF2 =25000;CF3=25000;CF4 =25000;CF =-25000;CPT IRR =-17.4%

The project should not be accepted because NPV is negative